OpenStudy (needhelpp101):

Solve the system using matrices. 3x + 4y = 18 3x + 4y = 18 A. (-2, 6) B. (-2, -6) C. (6, -2) D. No solution

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & 4 & 18\\ 3 & 4 & 18\end{matrix}\right]\] Using Gauss-Jordan elimination (we subtract first row from second, then simplify the first row so that the row's first entry is 1) we have:\[\left[\begin{matrix}1 & (3/4) & 6 \\ 0 & 0 & 0\end{matrix}\right]\] Since the reduced matrix has a row of zeroes, this indicated that there are infinitely many solutions. We can find all solutions by setting one variable equal to t and expressing the other in terms of that free variable. So let y = t, then x = 6 - (3/4)t,

OpenStudy (anonymous):

are you sure the second equation isn't different from the first?

OpenStudy (anonymous):

because if it's not, then we can only relate values of x to values of y that will satisfy that equation, in which case there will be infinitely many solutions

OpenStudy (needhelpp101):

I am sure....Solve the system using matrices. 3x + 4y = 18 3x + 4y = 18 A. (-2, 6) B. (-2, -6) C. (6, -2) D. No solution

OpenStudy (anonymous):

 SO ANSWER IS( -2,6)